OpenStudy (cakt07):

What is 20/400 written as a percent. Rounded to the Nearest tenth of a percent?

OpenStudy (cakt07):

@TuringTest can you help with this question?

OpenStudy (texaschic101):

divide it, and what is the decimal you get ?

OpenStudy (cakt07):

0.05

OpenStudy (cakt07):

Wait, then i round it?

OpenStudy (texaschic101):

now, to turn a decimal into a percent, move the decimal over two spaces to the right, or, you can just multiply it by 100.

OpenStudy (cakt07):

Okay i got 5

OpenStudy (texaschic101):

correct....5%

OpenStudy (cakt07):

OHHHHHHHH

OpenStudy (hanner_b_nanner):

you can use a proportion, too :)) 20/400=x/100...20x100=2000/400=5!

OpenStudy (cakt07):

Oh, Hold on i have to write this down in my notebook. So just wait k?

OpenStudy (texaschic101):

thats true...a proportion will also work

OpenStudy (hanner_b_nanner):

I find it easier that way, but that's just me! :)

OpenStudy (cakt07):

Ok what about 9/36

OpenStudy (cakt07):

And i can round it to the nearest tenth if neccassary.

OpenStudy (hanner_b_nanner):

9/36=x/100... 900/36=25

OpenStudy (texaschic101):

It might take a little longer to do it the way I do it. I usually turn the fraction to a decimal and then the decimal to a percent. 9 divided by 36 = n n x 100 = your answer as a percent

OpenStudy (cakt07):

O_O

OpenStudy (cakt07):

Okay well i got 25%

OpenStudy (texaschic101):

correct :)

OpenStudy (cakt07):